OpenStudy (anonymous):

Sum of the digits of a two digit number is 9. When we interchange the digits it is found that the resulting new number is greater than the original number by 27. What is the two-digit number?

OpenStudy (anonymous):

ur wrong

OpenStudy (anonymous):

36 3 + 6 = 9 36 + 27 = 63

OpenStudy (anonymous):

this is the answer Let the digits at tens place and ones place be x and 9 − x respectively. Therefore, original number = 10x + (9 − x) = 9x + 9 On interchanging the digits, the digits at ones place and tens place will be x and 9 − x respectively. Therefore, new number after interchanging the digits = 10(9 − x) + x = 90 − 10x + x = 90 − 9x According to the given question, New number = Original number + 27 90 − 9x = 9x + 9 + 27 90 − 9x = 9x + 36 Transposing 9x to R.H.S and 36 to L.H.S, we obtain 90 − 36 = 18x 54 = 18x Dividing both sides by 18, we obtain 3 = x and 9 − x = 6 Hence, the digits at tens place and ones place of the number are 3 and 6 respectively. Therefore, the two-digit number is 9x + 9 = 9 × 3 + 9 = 36